OpenStudy (anonymous):

Can somebody show me how to solve these? 3x+17y+4z=119 2x+3y+15z=87 10x+4y+7z=108

OpenStudy (precal):

Use elimination. \[-2(3x+17y+4z=119)\] \[-6x-34y-8z=-238\] \[3(2x+3y+15z=87)\] \[6x+9y+45z=261\] therefore -6x-34y-8z=-238 6x+9y+45z=261 when you add them become -25y+37z=23 do the same thing to the last 2 equations to get rid of the x variable. multiply a -5 to the second equation. Find y or z and then solve for x last.